OpenStudy (bossimbacon):

What is the solution to the system of equations represented by the two equations? y=2/3x y=-2/3x+4 ( , ) @robtobey

OpenStudy (bossimbacon):

@ParthKohli

OpenStudy (kropot72):

\[\large y=\frac{2}{3}x\ ..................(1)\] \[\large y=-\frac{2}{3}x+4\ ........(2)\] To solve for x, you need to plug the value for y given in equation (1) into equation (2).

OpenStudy (bossimbacon):

idk can you tell me the answer and then explain how you got it?

OpenStudy (kropot72):

Plugging in as described gives: \[\large \frac{2}{3}x=-\frac{2}{3}x+4\ .........(3)\] To solve equation (3) for x, the first step is to add (2/3)x to both sides of equation (3). Can you do that?

OpenStudy (kropot72):

To help a little more, I will set out the step as follows: \[\large \frac{2}{3}x+\frac{2}{3}x=(\frac{2}{3}x-\frac{2}{3}x)+4\ ........(4)\]

OpenStudy (bossimbacon):

:( i have to finish this soon. ughhh

OpenStudy (kropot72):

Solving (4) can be done in your head.

OpenStudy (bossimbacon):

@kittiwitti1 can you tell me the answer i have to finish this soon

OpenStudy (bossimbacon):

i will fan and medal

OpenStudy (kittiwitti1):

\[\text{Don't bribe people into helping you.}\]

OpenStudy (kittiwitti1):

No, I don't know the answer. I'm still looking through the posts.

OpenStudy (bossimbacon):

welp i honestly don't kno what to do, i don't understand this and i have to finish this soon soo

OpenStudy (kropot72):

@Bossimbacon I have explained exactly what to do. Are you prepared to do any work on this at all?

OpenStudy (kittiwitti1):

@kropot72 \[\text{It seems your explanation is too complex for him.}\]

OpenStudy (bossimbacon):

@kittiwitti1 if you're hear to "insult" me, then you can leave.

OpenStudy (bossimbacon):

here*

OpenStudy (kittiwitti1):

I'm not here* to "insult" you, I'm telling you to quit testing my patience when I'm trying to freaking HELP YOU. Unless your entire point was to annoy me. It's plausible.

OpenStudy (kittiwitti1):

...I'm still writing out the whole thing. You can go do something else while you wait.

OpenStudy (bossimbacon):

thank you for helping i appreciate it.

OpenStudy (kittiwitti1):

Well, whatever he tried to explain to you, I can't understand it either... and I tried to solve by making the equations equal each other, but that just deleted the x off the map entirely...

OpenStudy (kittiwitti1):

\[\text{WAIT I JUST HAD AN IDEA}!\]\[y = \frac{ 2 }{ 3 }x\]\[y= - \frac{2}{3} + 4\]Isolate x from the first equation\[x= \frac{3}{2}y\]and insert that x-value into the second equation. Does this help at all?

OpenStudy (kittiwitti1):

WHOOPS did it wrong! I solved and got -4 = 0 which is obviously wrong xD Use the idea that @kropot72 gave you:\[\frac{2}{3}x=- \frac{2}{3}x + 4\]\[\text{Now move all x-coefficients to the left side:}\]\[\frac{2}{3}x+\frac{2}{3}x=4\]\[\text{this becomes}\]\[\frac{4}{3}x=4\]\[\text{now isolate the x}\]\[x=4 \div \frac{4}{3}~becomes~x=4 \times \frac{4}{3}\]\[x= \frac{16}{3}\]\[\text{Did this help?}\]